Colter Company prepares monthly cash budgets. Relevant data from operating budgets for 2017 are as follows: anuary February $360,000 120,000 90,000 70,000 79,000 $400,000 125,000 100,000 75,000 85,000 Sales Direct materials purchases Direct labor Manufacturing overhead Selling and administrative expenses All sales are on account. Collections are expected to be 50% in the month of sale, 30% in the first month following the sale, and 20% in the second month following the sale. Sixty percent (60%) of direct materials purchases are paid in cash in the month of purchase, and the balance due is paid in the month following the purchase. All other items above are paid in the month incurred except for selling and administrative expenses that include $1,000 of depreciation per month Other data: 1. 2. 3. Credit sales: November 2016, $250,000; December 2016, $320,000 Purchases of direct materials: December 2016, $100,000 Other receipts: January-Collection of December 31, 2016, notes receivable $15,000; February-Proceeds from sale of securities $6,000 4. Other disbursements: February-Payment of $6,000 cash dividend The companys cash balance on January 1, 2017, is expected to be $60,000. The company wants to maintain a minimum cash balance of $50,000
Answer #1

Prepare schedule expected cash collection from customer :

January February
November 250000*20%=50000
December 320000*30%=96000 320000*20%=64000
January 360000*50%=180000 360000*30%=108000
February 400000*50%=200000
Total collections 326000 372000

Calculate expected payments for direct materials :

January February
December 100000*40%=40000
January 120000*60%=72000 120000*40%=48000
February 125000*60%=75000
Total payments 112000 123000